OpenStudy (anonymous):

square root(2x+15)-x=6 Can someone please help?! This is a quadratic I believe

OpenStudy (anonymous):

what are your directions?

OpenStudy (anonymous):

all it says is to solve the equation. it's college algebra and we're talking about quadratics

OpenStudy (anonymous):

get the square root on the side by itself. square both sides to remove the square root sign

OpenStudy (stamp):

\[(2x+15)^{1/2}-x=6\]\[(2x+15)^{1/2}=6+x\]\[2x+15=(6+x)^2\]Go from here.

OpenStudy (anonymous):



OpenStudy (anonymous):

foil out the right side then get all terms on the same side (set equation to 0)

OpenStudy (anonymous):

Thank you so much! :)